Write an equation of the line passing through (-3,5) and (2,6). Please give the answer in standard form.
The equation of the line in standard form is

Answer by Alan3354(69443)

Write an equation of the line passing through (-3,5) and (2,6). Please give the answer in standard form.
----------------
Find the slope, m
m = diffy/diffx = 1/5
-----
y - y1 = m*(x - x1) where (x1,y1) is either point.
y-6 = (1/5)*(x-2)
---
Change to the form you need.